OpenStudy (anonymous):

Two numbers differ by 9. The sum of their squares is 653. Find the numbers.

OpenStudy (anonymous):

Please help me.

OpenStudy (raden):

x-y=9 ---> x=y+9 x^2 + y^2 = 653 (y+9)^2 + y^2 = 653 2y^2 + 18y + 81 -653 = 0 2y^2 + 18y + 81 -572 = 0 y^2 + 9y - 286 = 0 (y+22)(y-13)=0 y=13 x=y+9=13+9=22

OpenStudy (anonymous):

Thanks! Such a big help.

OpenStudy (raden):

welcome

OpenStudy (raden):

wait, each numbers must be positive or may a negative numbers ?

OpenStudy (anonymous):

They are all positive.

OpenStudy (raden):

ok, it think that's all cleared

OpenStudy (anonymous):

Can you help me again this one is pretty hard for me. The length of a rectangle is 3cm longer than twice its width. The area is 119 square centimeters. What are the dimensions of the rectangle?

OpenStudy (raden):

l = 2w+3 A = 119 l*w = 119 (2w+3)w =119 2w^2 +3w -119 = 0 (2w + 17)(w - 7) = 0 w = 7 cm so, l = 2w+3 = 2(7)+3 = 17

OpenStudy (raden):

the dimensions of the rectangle : 17 x 7

OpenStudy (anonymous):

Genius! Thanks again!

OpenStudy (raden):

but, sometimes :p